OpenStudy (anonymous):

How do I solve this? 1/9(9y+18)-7=-1/5(10y-20)

OpenStudy (anonymous):

9/9(y + 18) - 7 = -10/5(y - 2); Then (y + 2) - 7 = -2 (y - 2). Isolate "y" and you will get y = 3.